OpenStudy (jianenriquez):

Where is the curve x^2y - xy -x^2 + 10x -12y -21 =0 symmetric?

OpenStudy (jianenriquez):

\[x^2y-xy-x^2-10x-12y-21=0\]

OpenStudy (triciaal):

can you get in the form ax^2 + bx + c = 0 ?

OpenStudy (jianenriquez):

erm i couldnt because of the x^2y

OpenStudy (triciaal):

-b/2a = (y + 10)/2(y-1)

OpenStudy (jianenriquez):

how did you get that?

OpenStudy (triciaal):

@jim_thompson5910 hi would you agree this would be the x-value for the vertex of a parabola or need another approach explicit

OpenStudy (jianenriquez):

i tried to see if i could get it into the general equation of the conic, i couldnt quite get it

jimthompson5910 (jim_thompson5910):

This is the graph that I get https://www.desmos.com/calculator/eguruel3b9 and I don't see any symmetry anywhere. Maybe there's a typo?

jimthompson5910 (jim_thompson5910):

I'm thinking that it should be x^2 - x*y -y^2 + 10x -12y -21 =0 if that's the case, then we get this graph https://www.desmos.com/calculator/ka8axdps59 which appears to have symmetry

OpenStudy (triciaal):

the original has 10x but below you have -10x which is correct?

jimthompson5910 (jim_thompson5910):

oh I missed that entirely. Thanks here is the fixed graph https://www.desmos.com/calculator/cvfq2jwvjy

OpenStudy (triciaal):

@jim_thompson5910 thanks for responding I see the graph but what is the solution? what would be the equation of the line of symmetry?

OpenStudy (jianenriquez):

Im currently reviewing for my Exit exam and i came across this question, it actually had choices a) x-axis b) y-axis c) origin d) none at all

OpenStudy (triciaal):

I see a line of reflection not of symmetry or would that be the same?

OpenStudy (triciaal):

multiple guess I would choose none sorry i could not be of more help

jimthompson5910 (jim_thompson5910):

there may be symmetry but I'm not 100% sure. There's definitely no x axis symmetry, no y axis symmetry, and there is no origin symmetry

jimthompson5910 (jim_thompson5910):

x axis symmetry happens when you can fold the graph over the x axis to have it line up with itself perfectly http://www.algebra-online.com/tutorials-4/radical-inequalities/articles_imgs/2199/symmet7.gif --------------------------------------------------------- y axis symmetry happens when you can fold the graph over the y axis to have it line up with itself perfectly http://mathbitsnotebook.com/Algebra1/FunctionGraphs/symgraph1.gif -------------------------------------------------------- Origin symmetry is where you can rotate a graph 180 degrees around the origin and have it line up with itself perfectly http://img.sparknotes.com/content/testprep/bookimgs/sat2/math2c/0054/odd.gif

jimthompson5910 (jim_thompson5910):

your graph meets none of those criteria so I'd go with `none` as the answer

OpenStudy (jianenriquez):

alright, thanks a lot :D